# Locality Alignment Improves Vision-Language Models

### v1

@@ B.1 Ablation study
We report the full results from our MaskEmbed ablation study in Table 4. These results inform our settings for the reconstruction target, data augmentations, mask sampling approach, training dataset and training duration. Separately, we also found in our early experiments that cosine similarity loss yielded similar results to MSE loss, and that varying the decoder depth and width did not lead to clear improvements; all our reported results therefore use a two-layer decoder with the same model dimension and number of attention heads as the pre-trained ViT. We describe each ablation parameter in detail below.
@@
Data augmentation. The least amount of data augmentation we can apply during MaskEmbed is a random crop and resize to the ViT’s resolution, in this case $224\times 224$ for both IN1k ViT-B and CLIP ViT-B. In addition to the random crop, we consider applying Mixup (Zhang et al., 2017), CutMix (Yun et al., 2019) and an AutoAugment recipe (Cubuk et al., 2018) as stronger augmentations. We find that Mixup and CutMix can help boost local probing performance but tend to hurt global probing performance. We opt to use the simple random crop in our remaining experiments, and we reason that strong augmentations are unnecessary because our masking leads to training each image with different targets in each epoch.
@@
(e) CLIP ViT-B/16 training data.
Table 4: MaskEmbed ablation study. We ablate several task design choices using our probing benchmark, including the teacher reconstruction target, data augmentations applied on top of masking, the mask sampling approach, and the training data for two pre-trained models (IN1k ViT-B/16 and CLIP ViT-B/16). We report the local and global probing performance for all runs. The teacher model results are written in gray, our default settings are highlighted in gray, and the best results are bolded.
@@ B.2 Additional results
Finally, Table 5 shows the results of running MaskEmbed on our full suite of pre-trained models. We observe that locality alignment improves local probing performance for all models trained with image-level supervision, and in most cases it also improves their global probing performance. The results are mixed for models trained with dense supervision: MAE, DINO and DINOv2 barely benefit from locality alignment (He et al., 2022, Caron et al., 2021, Oquab et al., 2023), and although BEiT and BEiTv2 do (Bao et al., 2021, Peng et al., 2022) this could be because we use checkpoints that are fine-tuned for IN1k classification. [^3] We use checkpoints available on timm at https://github.com/huggingface/pytorch-image-models. We also note that results between different models are sometimes not comparable due to differences in resolution and patch size. Surprisingly, DINOv2 is the best-performing model overall despite being a relatively weak backbone for VLMs (Karamcheti et al., 2024, Tong et al., 2024a); we interpret this to mean that DINOv2 is exceptionally good at detecting and localizing the set of classes in MSCOCO, which are relatively narrow and perhaps not indicative of the diverse images handled by VLMs.
@@ B.3 CLIPSelf comparison
Appendix C MaskEmbed Training Details
@@
Encoder. The encoder ViT is initialized from the pre-trained model weights and updated throughout training. Its input is an unmasked image, and its output is a sequence of patch embeddings that go through an additional linear output head. We experimented with re-initializing the final transformer block because these parameters are typically pre-trained only to pass information to the [CLS] token (Dosovitskiy et al., 2020, Radford et al., 2021), but this did not improve performance.
@@
Training instability. We encountered training instabilities in certain experiments, specifically a slow loss divergence that occurs partway through training. This type of instability has been reported in the literature with ViTs, with some works attributing it to saturation of the attention logits resulting in one-hot softmaxes (Zhai et al., 2023a); empirically, we were able to verify that diverged runs had a long tail of large attention logits. One common fix, QK-norm (Dehghani et al., 2023a, Chameleon Team, 2024), cannot be applied here because we fine-tune models that were pre-trained without QK-norm. We therefore use another approach that can be applied with a pre-trained model: logit soft-capping, where we use a tanh activation to constrain attention logits within a fixed range (Gemma Team et al., 2024). We adopt this approach in most of our MaskEmbed runs, including all runs that were used for training VLMs. We also had some success with increasing AdamW’s $\epsilon$ parameter and increasing the weight decay to 0.1, but these sometimes led to slower optimization.
@@
Hyperparameters. We report hyperparameters for our main MaskEmbed runs in Table 6. All models are trained with AdamW (Loshchilov and Hutter, 2017), slightly lower $\beta_{2}$ than the default value, moderate weight decay, minimal augmentations, gradient clipping, cosine learning rate schedule, and batch size 1024. All MaskEmbed runs are performed on a single node with 4 NVIDIA A100 SXM4 80GB GPUs.
@@ C.1 Additional perspectives
Relationship to masked image modeling. MaskEmbed bears some similarity to BERT-style masked imaging modeling (MIM) methods like MAE, MaskFeat and BEiT (He et al., 2022, Wei et al., 2022, Bao et al., 2021), but there are several important differences. 1) When encoding images, MIM methods mask the image at the input layer; MaskEmbed encodes the entire image and masks only at the output embedding layer. 2) MIM methods adopt static labels for each patch (although they typically only train on masked patches); we do not require labels for each patch embedding, and instead supervise predictions via their ability to reconstruct arbitrary masked teacher views. 3) Most MIM methods are designed for pre-training; MaskEmbed is a post-training method that can be applied to any pre-trained ViT backbone, including strong pre-training approaches that MIM methods struggle to match (e.g., CLIP, SigLIP; Radford et al. 2021, Zhai et al. 2023b).
Relationship to feature attribution. As described in the main text, our reconstruction objective in Equation 1 generalizes an existing feature attribution approach (Jethani et al., 2021, Covert et al., 2022). Given masked outputs $f(m(x))\in\mathbb{R}^{d}$ and a learned patch embedding model $g_{\theta}(x)\in\mathbb{R}^{n\times d}$, we can train the model to approximate $m^{\top}g_{\theta}(x)\approx f(m(x))$ for all $m$ using the following objective:
@@
Objective degeneracy. One potential concern about our approach is that the objective in Equation 1 is degenerate: it contains a trivial solution where the encoder acts as an identity function and the decoder replicates the teacher model, or $g_{\theta}(\cdot)=I(\cdot)$ and $h_{\phi}(\cdot)=f(\cdot)$. This solution is undesirable because it fails to encode rich semantics in each patch embedding, and when training a VLM it is equivalent to passing raw patch projections (similar to the Fuyu architecture; Bavishi et al. 2023). Given the strong performance we observe in practice from MaskEmbed, we reason that the trivial solution is avoided due to 1) the encoder’s strong initialization, and 2) the decoder’s small number of parameters and weak initialization. We tried training the encoder from scratch in our early experiments, and we found that it was important to use a shallow decoder to avoid simply preserving information with the encoder and offloading computation. However, the objective degeneracy does not seem like an issue when fine-tuning.
@@
Appendix D VLM Experiment Details & Additional Results
@@
Training data mixture. The Llava-1.5 training data mixture (Liu et al., 2024) consists of data sourced from several pre-existing datasets. These include synthetic instruction completions from the original Llava work (Liu et al., 2023c), a collection of existing VQA datasets (VQAv2, GQA, OCR-VQA, OK-VQA, A-OKVQA; Goyal et al. 2017, Hudson and Manning 2019, Marino et al. 2019, Mishra et al. 2019, Schwenk et al. 2022), captioning data (TextCaps; Sidorov et al. 2020), referring expression data (RefCOCO, Visual Genome; Kazemzadeh et al. 2014, Yu et al. 2016, Krishna et al. 2017), and ShareGPT data sourced from user conversations (ShareGPT, 2023). Our extended data mixture also includes the recent LVIS-Instruct-4V (Wang et al., 2023a) and LRV-Instruct (Liu et al., 2023b) datasets, which roughly double the number of training examples.
@@
Benchmark # Examples Prompt Type Scoring Details
VizWiz 4319 Open-ended VQA Some questions are unanswerable
VQAv2 214354 Open-ended VQA
GQA 12578 Open-ended Exact match
TextVQA (ocr) 5000 Open-ended VQA Prompt includes OCR dump
TextVQA (pure) 5000 Open-ended VQA No OCR dump
AI2D 15501 Multiple choice (4) Exact match
RefCOCO 10834 Bounding box Acc @ 0.5 IoU Spatial terms allowed
RefCOCO+ 10758 Bounding box Acc @ 0.5 IoU No spatial terms allowed
RefCOCOg 4896 Bounding box Acc @ 0.5 IoU Long object descriptions
OCID-Ref 18342 Bounding box Acc @ 0.25 IoU
VSR 1222 True/false Exact match
TallyQA (complex) 15598 Multiple choice (16) Exact match Involve filtering criteria
TallyQA (simple) 22991 Multiple choice (16) Exact match No filtering criteria
POPE 9000 Open-ended Exact match
D.1 Additional results
@@
First, Figure 11 shows a series of ablations for VLMs trained using different vision-language adapters. In the main text, we report that using the standard MLP adapter for aligned backbones degrades performance (see “Aligned MLP” vs. “Baseline MLP”) but that using the decoder as an adapter improves performance (see “Aligned Decoder”). To be sure that our improvements are due to locality alignment and not only the stronger adapter, we run several experiments using different adapter approaches for the baseline ViTs. First, we try training a transformer adapter from random initialization with the same size as the aligned model’s decoder; we find that this hurts performance compared to the MLP adapter (see “Baseline Transformer”), and we suspect that this is due to our VLM setup having insufficient training data to learn this module from random initialization. Previous works that successfully use transformer-based adapters have significantly more training data (Bai et al., 2023, Laurençon et al., 2024), so this result suggests that the decoder adapter is effective in part because it is initialized from pre-trained parameters.
@@
Finally, Figure 13 shows results from our feature fusion runs with DINOv2 (Oquab et al., 2023, Darcet et al., 2023). Our implementation of feature fusion follows Karamcheti et al. (2024): we concatenate the two output sequences along their embedding dimension and then pass this through a MLP adapter. As we describe in the main text, the fused backbones often lead to larger gains in core localization tasks, likely due to DINOv2’s excellent performance at dense prediction tasks (Oquab et al., 2023); however, it also leads the model to degrade in other ways, notably in VizWiz and TextVQA, which does not occur with our locality-aligned backbones.

### v2

@@ B.1 Ablation study
We report the full results from our MaskEmbed ablation study in Table 4. These results inform our settings for the reconstruction target, data augmentations, mask sampling approach, loss function, training dataset and training duration. Separately, we also found in our early experiments that varying the decoder depth and width did not lead to clear improvements; all our reported results therefore use a two-layer decoder with the same model dimension and number of attention heads as the pre-trained ViT. We describe each ablation parameter in detail below.
@@
Data augmentation. The minimum amount of data augmentation we can apply during MaskEmbed is a random crop and resize to the ViT’s resolution, in this case $224\times 224$ for both IN1k ViT-B and CLIP ViT-B. In addition to the random crop, we consider applying Mixup (Zhang et al., 2017), CutMix (Yun et al., 2019) and an AutoAugment recipe (Cubuk et al., 2018) as stronger augmentations. We find that Mixup and CutMix can help boost local probing performance but tend to hurt global probing performance. We opt to use the simple random crop in our remaining experiments, and we reason that strong augmentations are unnecessary because our masking leads to training each image with different reconstruction targets in each epoch.
@@
Loss function. We compare several reformulations of our loss function presented in Equation 1. Our base setting is the MSE reconstruction loss calculated over all patches, and we find that this performs slightly better than either a cosine similarity loss or a $\ell_{1}$ loss that penalize deviations differently. We also compare to reconstructing only the masked patches or only the unmasked patches; while the latter performs slightly better for global probing, we find that the best approach for both local and global probing is to simply reconstruct all patches, which differs slightly from works like MAE and BEiT (He et al., 2022; Bao et al., 2021). We reason that this is because all the patch embeddings are non-trivial reconstruction targets in our setup, compared to MAE where unmasked patches can be reconstructed with the identity function.
@@
local global
in1k teacher 43.50 51.04
cosine 45.55 51.37
$\ell_{1}$ 45.26 51.10
mse 45.66 51.43
mse (masked) 42.48 45.39
mse (unmasked) 45.00 51.67
(d) Loss function.
@@
(f) CLIP ViT-B/16 training data.
Table 4: MaskEmbed ablation study. We ablate several task design choices using our probing benchmark, including the teacher reconstruction target, data augmentations applied on top of masking, the mask sampling approach, loss function, and the training data for two pre-trained models (IN1k ViT-B/16 and CLIP ViT-B/16). We report the local and global probing performance for all runs. The teacher model results are written in gray, our default settings are highlighted in gray, and the best results are bolded.
@@ B.2 Additional results
Finally, Table 5 shows the results of running MaskEmbed on our full suite of pre-trained models. We observe that locality alignment improves local probing performance for all models trained with image-level supervision, and in most cases it also improves their global probing performance. The results are mixed for models trained with dense supervision: MAE, DINO and DINOv2 barely benefit from locality alignment (He et al., 2022; Caron et al., 2021; Oquab et al., 2023), and although BEiT and BEiTv2 do (Bao et al., 2021; Peng et al., 2022) this could be because we use checkpoints that are fine-tuned for IN1k classification. [^4] We use checkpoints available on timm at https://github.com/huggingface/pytorch-image-models. We also note that results between different models are sometimes not comparable due to differences in resolution and patch size. Surprisingly, DINOv2 is the best-performing model overall despite being a relatively weak backbone for VLMs (Karamcheti et al., 2024; Tong et al., 2024a); we interpret this to mean that DINOv2 is exceptionally good at detecting and localizing the set of classes in MSCOCO, which are relatively narrow and perhaps not indicative of the diverse images handled by VLMs.
@@ B.3 CLIPSelf comparison
B.4 Masked image modeling comparisons
In Section 3.3, we briefly discuss how locality alignment only works when the teacher model appropriately handles masked images, and how this can be encouraged by fine-tuning with random masking (Frye et al., 2020; Covert et al., 2021; Jain et al., 2022) before using locality alignment. We now explore the impact of including this additional fine-tuning stage.
In particular, we explore fine-tuning the ViT network with randomly masked input patches, and an objective of reconstructing embeddings from a frozen version of itself that processes entire images. Intuitively, this teaches the model to predict the semantics of the entire image by making its best guess for the missing patch contents, which is ideal behavior for the teacher model in locality alignment. We remark that this is similar to masked image modeling methods like MAE (He et al., 2022), but performing this task with rich features is known to work better than with raw pixels as the reconstruction target (Wei et al., 2022), and is the basis of recent masked image modeling works like BEiT v2 and EVA.
In Table 6, we test using locality alignment in different combinations with masked image modeling. Similar to locality alignment, we train for 5 epochs with IN21k and set other hyperparameters identically. We find that applying either fine-tuning approach to the original CLIP ViT-B backbone improves local and global probing, but that the gains are significantly higher for locality alignment. The gains on local probing are further improved if we perform masked image modeling followed by locality alignment, as predicted, which suggests that masked image modeling is a powerful precursor for locality alignment. Note that it could in principle be up-streamed into pre-training by simply dropping random patches, similar to FLIP (Li et al., 2023b). The best overall local probing performance is achieved by applying a subsequent round of masked image modeling fine-tuning, but this significantly degrades global probing. Meanwhile, the best overall global probing performance is achieved by applying locality alignment directly on the original CLIP ViT-B. For a relatively simple procedure that performs well at both probing tasks, a single round of masked image modeling followed by locality alignment is competitive at both evaluations, but we did not use this in any other experiments.
Table 6: Combining masked image modeling with locality alignment. We compare local and global probing performance for CLIP ViT-B models with different sequences of masked image modeling (MIM) and locality alignment. We find the the local probing performance can be significantly improved by performing locality alignment after an initial masked image modeling stage.
Model Teacher local global
CLIP ViT-B 44.63 52.61
CLIP ViT-B (MIM) CLIP ViT-B 45.80 52.98
CLIP ViT-B (Align) CLIP ViT-B 46.32 54.55
CLIP ViT-B (MIM/Align) CLIP ViT-B (MIM) 47.30 53.63
CLIP ViT-B (MIM/Align/MIM) CLIP ViT-B (MIM/Align) 48.54 51.05
CLIP ViT-B (MIM/Align/MIM/Align) CLIP ViT-B (MIM/Align/MIM) 47.80 53.34
Appendix C ImageNet Classification
Our vision-centric experiments in Section 4 rely on the probing benchmark (see Appendix B), which assesses both local and global feature extraction but with a relatively narrow set of classes. To further test global semantic understanding and verify that it does not degrade with locality alignment, we also consider IN1k classification. We adopt an end-to-end fine-tuning setup with hyperparameters similar to those in OpenCLIP (Cherti et al., 2023), and we focus on CLIP backbones only for simplicity (Radford et al., 2021). One difference in our setup is that we replace the standard linear head for a transformer output head, because locality alignment teaches the model to output relevant semantics in all the output embeddings rather than just the [CLS] token.
The results are shown in Table 7, and show that the classification accuracy does not degrade but instead improves after locality alignment, across three ViT architecture variants. These results echo those in Table 5 for global probing, but in this case with end-to-end fine-tuning rather than a frozen backbone, and also a more challenging classification task. We attribute the improved performance to 1) the use of a full mask in MaskEmbed that leads to preserving global image understanding, and 2) an overall more challenging task that leads to stronger and more steerable internal features.
Table 7: IN1k classification accuracy. For each model, we perform end-to-end fine-tuning for 50 epochs before and after locality alignment, and we report the top-1 accuracy.
Model Baseline Aligned
CLIP ViT-B 82.6 83.1
CLIP ViT-L 85.9 86.3
CLIP ViT-L @ 336px 86.4 87.0
Appendix D MaskEmbed Training Details
@@
Encoder. The encoder ViT is initialized from the pre-trained model weights and updated throughout training. Its input is an unmasked image, and its output is a sequence of patch embeddings that go through an additional linear output head. We experimented with re-initializing the final transformer block because these parameters are typically pre-trained only to pass information to the [CLS] token (Dosovitskiy et al., 2020; Radford et al., 2021), but this did not improve performance.
@@
Training instability. We encountered training instabilities in certain experiments, specifically a slow loss divergence that occurs partway through training. This type of instability has been reported in the literature with ViTs, with some works attributing it to saturation of the attention logits resulting in one-hot softmaxes (Zhai et al., 2023a); empirically, we were able to verify that diverged runs had a long tail of large attention logits. One common fix, QK-norm (Dehghani et al., 2023a; Chameleon Team, 2024), cannot be applied here because we fine-tune models that were pre-trained without QK-norm. We therefore use another approach that can be applied with a pre-trained model: logit soft-capping, where we use a tanh activation to constrain attention logits within a fixed range (Gemma Team et al., 2024). We adopt this approach in most of our MaskEmbed runs, including all runs that were used for training VLMs. We also had some success with increasing AdamW’s $\epsilon$ parameter and increasing the weight decay to 0.1, but these sometimes led to slower optimization.
@@
Hyperparameters. We report hyperparameters for our main MaskEmbed runs in Table 8. All models are trained with AdamW (Loshchilov & Hutter, 2017), slightly lower $\beta_{2}$ than the default value, moderate weight decay, minimal augmentations, gradient clipping, cosine learning rate schedule, and batch size 1024. All MaskEmbed runs are performed on a single node with 4 NVIDIA A100 SXM4 80GB GPUs.
@@ D.1 Additional perspectives
Relationship to masked image modeling. MaskEmbed bears some similarity to BERT-style masked imaging modeling (MIM) methods like MAE, MaskFeat and BEiT (He et al., 2022; Wei et al., 2022; Bao et al., 2021), but there are several important differences. 1) When encoding images, MIM methods mask the image at the input layer; MaskEmbed encodes the entire image and masks only at the output embedding layer. 2) MIM methods adopt static labels for each patch (although they typically only train on masked patches); we do not require labels for each patch embedding, and instead supervise predictions via their ability to reconstruct arbitrary masked teacher views. 3) Most MIM methods are designed for pre-training; MaskEmbed is a post-training method that can be applied to any pre-trained ViT backbone, including strong pre-training approaches that MIM methods struggle to match (e.g., CLIP, SigLIP; Radford et al. 2021; Zhai et al. 2023b).
Relationship to feature attribution. As described in the main text, our reconstruction objective in Equation 1 generalizes an existing feature attribution approach (Jethani et al., 2021; Covert et al., 2022). Given masked outputs $f(m(x))\in\mathbb{R}^{d}$ and a learned patch embedding model $g_{\theta}(x)\in\mathbb{R}^{n\times d}$, we can train the model to approximate $m^{\top}g_{\theta}(x)\approx f(m(x))$ for all $m$ using the following objective:
@@
Objective degeneracy. One potential concern about our approach is that the objective in Equation 1 is degenerate: it contains a trivial solution where the encoder acts as an identity function and the decoder replicates the teacher model, or $g_{\theta}(\cdot)=I(\cdot)$ and $h_{\phi}(\cdot)=f(\cdot)$. This solution is undesirable because it fails to encode rich semantics in each patch embedding, and when training a VLM it is equivalent to passing raw patch projections (similar to the Fuyu architecture; Bavishi et al. 2023). Given the strong performance we observe in practice from MaskEmbed, we reason that the trivial solution is avoided due to 1) the encoder’s strong initialization, and 2) the decoder’s small number of parameters and weak initialization. We tried training the encoder from scratch in our early experiments, and we found that it was important to use a shallow decoder to avoid simply preserving information with the encoder and offloading computation. However, the objective degeneracy does not appear to be an issue when fine-tuning.
@@
Appendix E VLM Experiment Details & Additional Results
@@
Training data mixture. The Llava-1.5 training data mixture (Liu et al., 2024) consists of data sourced from several pre-existing datasets. These include synthetic instruction completions from the original Llava work (Liu et al., 2023c), a collection of existing VQA datasets (VQAv2, GQA, OCR-VQA, OK-VQA, A-OKVQA; Goyal et al. 2017; Hudson & Manning 2019; Marino et al. 2019; Mishra et al. 2019; Schwenk et al. 2022), captioning data (TextCaps; Sidorov et al. 2020), referring expression data (RefCOCO, Visual Genome; Kazemzadeh et al. 2014; Yu et al. 2016; Krishna et al. 2017), and ShareGPT data sourced from user conversations (ShareGPT, 2023). Our extended data mixture also includes the recent LVIS-Instruct-4V (Wang et al., 2023a) and LRV-Instruct (Liu et al., 2023b) datasets, which roughly double the number of training examples.
@@
Benchmark # Examples Prompt Type Scoring Details VizWiz 4319 Open-ended VQA Some questions are unanswerable VQAv2 214354 Open-ended VQA GQA 12578 Open-ended Exact match TextVQA (ocr) 5000 Open-ended VQA Prompt includes OCR dump TextVQA (pure) 5000 Open-ended VQA No OCR dump AI2D 15501 Multiple choice (4) Exact match RefCOCO 10834 Bounding box Acc @ 0.5 IoU Spatial terms allowed RefCOCO+ 10758 Bounding box Acc @ 0.5 IoU No spatial terms allowed RefCOCOg 4896 Bounding box Acc @ 0.5 IoU Long object descriptions OCID-Ref 18342 Bounding box Acc @ 0.25 IoU VSR 1222 True/false Exact match TallyQA (complex) 15598 Multiple choice (16) Exact match Involve filtering criteria TallyQA (simple) 22991 Multiple choice (16) Exact match No filtering criteria POPE 9000 Open-ended Exact match
E.1 Additional results
@@
First, Figure 11 shows a series of ablations for VLMs trained using different vision-language adapters. In the main text, we report that using the standard MLP adapter for aligned backbones degrades performance (see “Aligned MLP” vs. “Baseline MLP”) but that using the decoder as an adapter improves performance (see “Aligned Decoder”). To be sure that our improvements are due to locality alignment and not only the stronger adapter, we run several experiments using different adapter approaches for the baseline ViTs. First, we try training a transformer adapter from random initialization with the same size as the aligned model’s decoder; we find that this hurts performance compared to the MLP adapter (see “Baseline Transformer”), and we suspect that this is due to our VLM setup having insufficient training data to learn this module from random initialization. Previous works that successfully use transformer-based adapters have significantly more training data (Bai et al., 2023; Laurençon et al., 2024), so this result suggests that the decoder adapter is effective in part because it is initialized from pre-trained parameters.
@@
Finally, Figure 13 and Figure 14 show results from our feature fusion runs with DINOv2 (Oquab et al., 2023; Darcet et al., 2023). Our implementation of feature fusion follows Karamcheti et al. (2024): we concatenate the two output sequences along their embedding dimension and then pass this through a MLP adapter. As we describe in the main text, the fused backbones often lead to larger gains in core localization tasks, likely due to DINOv2’s excellent performance at dense prediction tasks (Oquab et al., 2023); however, it also leads the model to degrade in other ways, notably in VizWiz and TextVQA, which does not occur with our locality-aligned backbones. Overall, the more robust improvements from locality alignment make it an appealing option to improve localization tasks without negatively impacting other model capabilities.
@@
Figure 14: VLM comparison with DINOv2 feature fusion. We compare VLMs with locality-aligned backbones to fusing features between CLIP/SigLIP and DINOv2. TextVQA benchmarks are not shown for CLIP ViT-L + DINOv2 fusion due to the accuracy lying outside the display range, more than three standard deviations below the mean performance.